I have a total of 35 coins consisting of dimes and nickels. THe value of these coins is $3.30. How many dimes and nickels do i have?
Answer by stanbon!(97)

Let the number of nickels be "x".
Then number of dimes is "35-x".
Value of nickels is (5)x cents.
Value of dimes is (10)(35-x)= (350 -10x) cents.
EQUATION:
Value of nickels + Value of dimes is 330 cents
5x +350-10x = 330
-5x = -20
x = 4 (number of nickels)
35-x = 31 (number of dimes)
Cheers,
Stan H.